a particular algorithm increases in time as the number of operations n increases.

Suppose the time complexity of this algorithm is given by:

f(n)=4n2+5n2*log(n)

Show that f(n) is O(g(n)) for g(n) = n3

1
Expert's answer
2021-11-30T13:26:50-0500

"f(n)=4n^2+5n^2*log(n)" , "g(n) = n^3"

Now to show "f(n)=O(g(n))" when number of operation increases i.e., "n\rightarrow \infty" then

"\frac{f(n)}{g(n)}=\frac{4n^3+5n^2.log\ n}{n^3}=4+\frac{5.log\ n}{n}"

Now, "\lim_{n\to\infty}" "\frac{f(n)}{g(n)}=4+\lim_{n\to\infty}\frac{5.log\ n}{n}"

"=4+5\lim_{n\to\infty} \frac{1/n}{1}" [Using L'Hospital Rule]

"=4+5\times0=4"

So, "|\frac{f(n)}{g(n)}|\rightarrow 4" as "n\rightarrow\infty"

So, "f(n)=O(g(n))"